# The Impact of COVID-19 on Co-authorship and Economics Scholars’ Productivity

### v1

@@ 5 Empirical Results
The estimated peer effect is significant for pre-Covid times (2018-2019), but not for Covid times (2020-2021). This implies that economic scholars exhibit conformity in the number of publications before the schools switch to virtual mode. However, there is a sign that while working at home, economic scholars are cooperating with a more diverse group of co-authors in terms of productivity, i.e. more collaborations among prolific scholars and scholars who publish less are expected. Recent productivity, represented by the scholar’s average number of publications during the previous 3 years, is an important predictor of a scholar’s productivity in the future. Scholars who are more prolific in the past 3 years would be more productive in the future. Compared to scholars whose total citations are below 2,000, scholars who have more citations are likely to publish more. Scholars’ years in academia significantly affect their productivity during Covid times. Emerging scholars produce relatively more than senior scholars. Gender plays a significant role only during the pre-Covid time.
Productivity is also related to the economics sub-fields that scholars work in. For example, econometricians publish relatively fewer papers in 2018-2019. Scholars who work in health economics are more prolific, and the effect is stronger during Covid times. The significant and positive Covid index implies that during 2019-2021, the higher proportion of a scholar’s papers is related to Covid, the more publications he or she would have. After controlling for the Covid index, the coefficient of scholars working in the field of health economics drops from $0.33$ to $0.30$.
For contextual effect, during Covid times, collaborating with macroeconomists and established scholars who have 100-2,000 citations help publish more papers.
@@ 6 Conclusion
In this study, we examine the influence of peer effects on the productivity of economics scholars in both pre-pandemic and pandemic periods. Findings reveal that scholars tend to coauthor with others with similar productivity during the pre-pandemic time, but this conformity is not found during the pandemic time. Productivity during the previous 3 years helps predict productivity in the near future. Citation count is positively correlated with scholars’ productivity at both times, while academic experience only affects the pandemic time. Female scholars are reported to publish less during the pre-pandemic time, but the effect vanishes during the pandemic time. This paper contributes to the growing body of literature on peer effects and academic productivity, and the insights help inform policies aimed at fostering research collaboration and enhancing productivity within the academic community.
While our research offers important insights, it is not without limitations. For example, in the present study, we assess an economics scholar’s productivity based on the number of publications. This may be questionable because factors like the quality of the research and the type of publications are not considered. More comprehensive measures could be used by including citation counts, journal impact factors, type of publications, altmetrics, etc. Relative to other data sources, although Google Scholar has several advantages, such as broader coverage of different types of publication, the information being up-to-date, etc., it may offer inaccurate author profiles and results of inconsistent accuracy [43]. Future research may benefit from incorporating additional data sources from journal websites or other online platforms to ensure a more comprehensive and accurate assessment of scholars’ productivity and co-authorship patterns. In addition, we only investigate the peer effect within a 4 years period. As the pandemic continues to affect the global academic community, it would be crucial to monitor the long-term implications of these changing collaboration patterns and their impact on research productivity. Further exploration of factors that facilitate or hinder research collaboration during such crises can help guide the development of effective strategies and policies to support and strengthen the academic community in times of unprecedented challenges.

### v2

@@ 5 Empirical Results
The estimated peer effect is significant for pre-Covid period from 2018 to 2019, but not significant for Covid times from 2020 to 2021. This estimation implies that economic scholars exhibited conformity in the number of publications before schools switched to virtual mode. However, there is an indication that while working from home, economic scholars collaborated with a more diverse group of co-authors; in terms of productivity, there were more collaborations among prolific scholars than those who publish less frequently.
Our results illustrate that scholars’ recent productivity, represented by their average number of publications over the previous three years, is an important predictor of future productivity. In other words, scholars who were more prolific in the past three years tend to be more productive in the future. We also find that compared to scholars with fewer than 2,000 total citations, those with more citations are likely to publish more. Scholars’ years in academia significantly affect their productivity during Covid times, with emerging scholars producing relatively more than senior scholars. In addition, gender plays a significant role only during the pre-Covid period.
Moreover, our results show productivity is also related to the economics sub-fields that scholars work in. For example, econometricians collaborated and published relatively fewer papers in 2018-2019. Scholars working in health economics are more prolific, and this effect is stronger during Covid times. The significant and positive Covid index implies that during 2019-2021, the higher the proportion of a scholar’s papers related to Covid, the more publications he or she would have. After controlling for the Covid index, the coefficient for scholars working in the field of health economics drops from $0.33$ to $0.30$.
Furthermore, regarding the contextual effect, during Covid times, collaborating with macroeconomists and established scholars who have 100-2,000 citations helped increase the number of publications.
@@ 6 Conclusion
In this study, we examine the influence of peer effects on the productivity of economic scholars in both pre-pandemic and during pandemic periods. The findings reveal that scholars tended to collaborate with others of similar level of productivity during the pre-pandemic period, but this conformity was not observed during the pandemic. Productivity during the previous three years helps predict productivity in the near future. Citation count is positively correlated with scholars’ productivity in both periods, while academic experience only affects productivity during the pandemic. Female scholars were reported to publish less during the pre-pandemic period, but this effect vanishes during the pandemic. This paper contributes to the growing body of literature on peer effects and academic productivity, offering insights that can help inform policies aimed at fostering research collaboration and enhancing productivity within the academic community.
While our research offers important insights, it is not without limitations. For example, in the present study, we assess each economics scholar’s productivity based on her/his number of publications. This measure may be questionable because it does not account for factors like the quality of the research and the type of publications. More comprehensive measures could include citation counts, journal impact factors, types of publications, altmetrics, etc. Although Google Scholar has several advantages, such as broader coverage of different types of publications and up-to-date information, it may offer inaccurate author profiles and results of inconsistent accuracy (Falagas et al., 2008). Future research may benefit from incorporating additional data sources from journal websites or other online platforms to ensure a more comprehensive and accurate assessment of scholars’ productivity and co-authorship patterns. Moreover, future research can use our findings to future explore the literature on Diversity, Equity, and Inclusion (DEI) in academia.
In addition, in our paper, we only investigate the peer effects within a four-year period. The Covid-19 pandemic has unprecedentedly affected the global academic community, potentially causing a paradigm shift as many scholars leave academia during and after the pandemic. Therefore, it is crucial to continue monitoring and studying the long-term implications of these collaboration patterns and their impact on research productivity. Further exploration of the factors that facilitate or hinder research collaboration during such crises can help guide the development of effective strategies and policies to support and strengthen the academic community in times of unprecedented challenges.